The ratio of the number of boys to the number of girls at school is 4:5. If there are 120 boys how many students are there altogether?
Answer by TimothyLamb(4379)

4/5 = 120/x
x = 120/(4/5)
x = 150 number of girls
---
total students:
120 + 150 = 270
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---